OpenStudy (anonymous):

-5.3> 6.7 + 4.3 + q -

OpenStudy (anonymous):

6.7+4.3=11. Subtract 11 from both sides to isolate q

jhonyy9 (jhonyy9):

-5,3 >=11 +q -5,3 -11>=q -16,3>=q